OpenStudy (anonymous):

Which line is perpendicular to the line y = 1/2x + 2 ?

OpenStudy (dmb1215):

y=-2x

OpenStudy (anonymous):

Slope = 1.000/2.000 = 0.500 2. x-intercept = 4/-1 = -4.00000 3. y-intercept = 4/2 = 2 My issue is that I'm not understanding how to get the accurate perpendicular line.

OpenStudy (dmb1215):

ok heres a hint for finding perpendicular lines the slope of the perpendulcar line is the negative reciprocal of the original

OpenStudy (dmb1215):

and you dont divide the slopes

OpenStudy (anonymous):

So what you are saying is: the negative reciprocal of 1/2x + 2 is -2/1x +2?

OpenStudy (dmb1215):

it wont be +2 because the +2 is the y intercept

OpenStudy (dmb1215):

for the original line

OpenStudy (dmb1215):

well i take that back it can be

OpenStudy (anonymous):

So if y=-2x, how do I calculate the perpendicular line? This is confusing to me.

OpenStudy (dmb1215):

y=1/2x+2 Y=-2x See what i did and what do you mean calculate

OpenStudy (anonymous):

Sorry, did not mean calculate, I meant find

OpenStudy (dmb1215):